OpenStudy (anonymous):

6x^4/4x^9 times 12x^2/3x^5

OpenStudy (anonymous):

start with \[\frac{6x^4\times 12x^2}{4x^9\times 3x^5}\] then add the exponents in the numerator and denominator and cancel

OpenStudy (anonymous):

72x^6/12x^14

OpenStudy (anonymous):

you get \[\frac{6\times 12\times x^{4+2}}{12\times x^{9+5}}=\frac{6\times 12\times x^6}{12\times x^{14}}\] \[\frac{6\times \cancel{12} x^6}{\cancel{12}x^{14}}=\frac{6x^6}{x^{14}}\] \[\frac{6}{x^{14-6}}=\frac{6}{x^8}\]